OpenStudy (anonymous):

Show all work. A disc jockey has 10 songs to play. Seven are slow songs, and three are fast songs. Each song is to be played only once. In how many ways can the disc jockey play the 10 songs if The songs can be played in any order. The first song must be a slow song and the last song must be a slow song. The first two songs must be fast songs.

OpenStudy (anonymous):

any order 10!

OpenStudy (anonymous):

oops i had slow and fast backwards.

OpenStudy (anonymous):

huh!

OpenStudy (anonymous):

i screwed up on #2 not i have to think.

OpenStudy (anonymous):

zarkon?

OpenStudy (anonymous):

7 for the first, 6 for the last, and 8*7*6*5*4*3 for the middle so i think it is 7*6/8*7*6*5*4*3

OpenStudy (anonymous):

How can the first song be both fast and slow?

OpenStudy (anonymous):

first two slow 3*2 then 8*7*6*5*4*3

OpenStudy (zarkon):

is there a reason you are dropping the 2?

OpenStudy (zarkon):

\[10!\] \[_7P_2\cdot 8!\] \[_3P_2\cdot 8!\]

OpenStudy (anonymous):

i have no clue what yall are doing please explain!

OpenStudy (zarkon):

\[_7P_2=7\times 6\] \[_3P_2=3\times 2\]

OpenStudy (zarkon):

1) 3628800 2) 1693440 3) 241920

OpenStudy (anonymous):

ok so it was right yes? i mean what i wrote

OpenStudy (anonymous):

oh no it was wrong. i didn't put the 2 at the end. doh

OpenStudy (anonymous):

first two slow 3*2 then 8*7*6*5*4*3*2 so 3*2*8*7*6*5*4*3*2

OpenStudy (anonymous):

i would like to think that was a typo

OpenStudy (zarkon):

I would too

OpenStudy (anonymous):

lol!